The sum of three numbers is 138. The second is 5 more than the smallest and the third is 10 more than the smallest. Find the numbers.
Answer by sachi(548)

let the smallest= x
the 2nd= x+5 & 3rd=x+10
nowx+x+5+x+10=138
or 3x=138-15=123
so x=41
so the nos are 41,46,51 ans